Write the differential equation y'e^x + ye^2x - six = 0 in the standard form


1
Expert's answer
2021-11-05T06:04:40-0400

Let us write the differential equation "y'e^x + ye^{2x} - \sin x = 0" in the standard form.

It follows that "y'e^x + ye^{2x} =\sin x," and hence "y' + e^{x}y =e^{-x}\sin x" is the standard form of the differential equation "y'e^x + ye^{2x} - \sin x = 0."